# RS-Agent: Automating Remote Sensing Tasks through Intelligent Agent

### v2

@@ 1 Introduction
By aligning visual and textual information, Multi-modal Large Language Models (MLLMs) [Radford et al., 2021, Sun et al., 2023] have made considerable progress across diverse remote sensing image interpretation tasks, including image captioning [Tao et al., 2023], visual question answering [Bashmal et al., 2023], and semantic scene understanding [Hu et al., 2025, Kuckreja et al., 2023], making remote sensing data more accessible to non-expert users. However, despite these advancements, MLLMs still face challenges in consistently handling multiple tasks across diverse remote sensing image modalities and resolutions. Firstly, remote sensing data encompass a wide range of modalities with varying resolutions. Consequently, general-purpose models lack the flexibility to quickly accommodate new modalities without computationally expensive retraining. In addition, most existing MLLMs are primarily optimized for instruction-following or descriptive tasks, such as image captioning and visual question answering. These generalist models typically do not provide expert-level responses comparable to those specialized models customized for specific tasks [Sun et al., 2021, Li et al., 2021a, Wang et al., 2022, Zhu et al., 2022]. These challenges emphasize the necessity for an expert-aware system capable of interacting with human users, and automatically utilizing specialized models to address the demands of real-world remote sensing applications.
It is believed that AI Agent [Xi et al., 2025] is a promising solution to these challenges. AI agents are autonomous systems capable of understanding user intent, planning multi-step tasks, invoking appropriate tools, and refining their decisions based on intermediate outcomes [Reed et al., 2022, Park et al., 2023, Wang et al., 2024d]. Recent advancements in agent frameworks—often powered by large language models—have demonstrated strong potential in coordinating multiple tools, decomposing complex tasks, and flexibly adapting to diverse user goals through modular design [Ruan et al., 2023, Xi et al., 2025]. These agents can bridge the gap between high-level semantic instructions and low-level tool execution operations. Despite growing interest, such agent-based approaches remain largely underexplored in the remote sensing domain. Although a few works in the remote sensing field use the term “agent” in their paper [Du et al., 2023, Liu et al., 2024, Zhu et al., 2024], they primarily focus on specific tasks and fail to propose a general-purpose agent. RS-ChatGPT [Guo et al., 2024a] integrates ChatGPT [Brown et al., 2020] with a collection of pre-trained remote sensing networks to perform tasks such as object detection, scene classification, and image captioning. However, RS-ChatGPT encounters limitations when scaling to a broader range of tools or addressing tasks that demand specialized domain knowledge.
@@
The core functionality of RS-Agent lies in its ability to accurately interpret user instructions and convert them into precise and executable workflows. To this end, we design a carefully structured agent architecture that integrates modular system prompts with a long-term memory mechanism. This design empowers the Central Controller to accurately interpret user intent and dynamically plan workflows by leveraging current inputs, tool outputs, and contextual memory. Robustness and adaptability are critical for such systems, especially when handling diverse and ambiguous user instructions. Recent studies have emphasized the importance of stable task-level decision-making under varying task distributions and execution contexts [Wang et al., 2024a, Lv et al., 2024]. In the Solution Space, to empower RS-Agent with the ability to reason and operate like a professional remote sensing analyst, we build a Solution Database that stores professional knowledge on how to use various remote sensing tools effectively. We further propose a Task-Aware Retrieval method to retrieve and comprehend expert knowledge, enabling accurate tool selection and step-by-step task decomposition tailored to the user’s query. To address the domain knowledge limitations of general-purpose models [Brown et al., 2020, Touvron et al., 2023, Bai et al., 2023], we develop a dedicated Knowledge Database, tailored to remote sensing targets. Building on this, we propose DualRAG, a novel retrieval-augmented generation method that enables the agent to retrieve more relevant and accurate domain-specific knowledge.
RS-Agent accurately interprets user intent and offers flexible integration of new tools to support emerging applications. We conduct extensive experiments across 9 datasets and 18 remote sensing tasks. Experiments reveal that the RS-Agent significantly enhances task planning accuracy—achieving over 95% accuracy, substantially outperforming the SOTA baseline RS-ChatGPT [Guo et al., 2024a]. Furthermore, by coordinating multiple tools, RS-Agent consistently surpasses single-model performance across diverse applications, including scene classification, remote sensing visual question answering, and object counting. In addition, the proposed DualRAG method markedly improves the quality of domain-specific knowledge retrieval. It is also worth noting that this work does not aim to develop new LLM architectures. Instead, RS-Agent is designed to be compatible with both open-source and proprietary LLMs. We validate the generality and adaptability of our framework through experiments with several state-of-the-art models, including ChatGPT [Brown et al., 2020], LLaMA [Touvron et al., 2023], Qwen [Yang et al., 2024], and DeepSeek [Guo et al., 2025], demonstrating the generality and adaptability of our framework.
@@ 2.1 Remote Sensing Multimodal Large Language Models
The integration of Multimodal Large Language Models (MLLMs) into remote sensing has opened new opportunities for human-AI interaction in earth observation tasks. MLLMs aim to bridge visual and semantic spaces, enabling the processing and interpretation of remote sensing imagery based on natural language inputs. Early applications of MLLMs in remote sensing primarily focused on a limited set of tasks, such as image captioning [Shi and Zou, 2017, Wang et al., 2020b] and visual question answering (VQA) [Lobry et al., 2020]. As research progressed, MLLMs began to support a broader range of vision-language tasks [Wang et al., 2025], each addressing specific challenges in remote sensing interpretation. RSGPT [Hu et al., 2025] integrates both image captioning and VQA. GeoChat [Kuckreja et al., 2023] extended this capability to visual geo-localization via interactive dialogue. LHRS-Bot [Muhtar et al., 2024] is a general-purpose MLLM built upon large-scale datasets and a novel multi-level vision-language alignment strategy, achieving state-of-the-art performance in classification, visual question answering, and visual grounding tasks. EarthVQANet [Wang et al., 2024c] introduces a novel multi-task framework combining semantic segmentation and VQA to enhance spatial reasoning in remote sensing imagery. SkySenseGPT [Luo et al., 2024] leveraged knowledge graphs to enhance relational reasoning and complex target understanding. Beyond single-image interpretation tasks, a number of MLLM-based models have also been developed to address the interpretation of multi-scene remote sensing data. ChangeCLIP [Dong et al., 2024] filled a critical gap by tackling change captioning between multi-temporal images. SkyEyeGPT [Zhan et al., 2025] introduced video captioning for remote sensing. UniRS [Li et al., 2024] further advanced the field by integrating single-frame, bi-temporal, and video inputs into a unified model designed specifically for temporal remote sensing analysis.
@@ 2.2 AI Agent
AI agents are autonomous systems designed to perceive their environment, process information, make decisions, and execute actions to achieve specific objectives [Wu et al., 2023]. Equipped with the ability to learn from interactions, adapt to new data, and continuously improve performance, these agents are capable of functioning effectively across a wide range of application domains [Park et al., 2023, Durante et al., 2024].
While AI agents have been widely explored in general domains, their application in the field of remote sensing remains limited. Only a few agent-based systems have been proposed, typically addressing some specific defined tasks. TreeGPT [Du et al., 2023], for instance, is designed for forestry applications, enabling individual tree segmentation and ecological parameter estimation from remote sensing imagery. Change-Agent [Liu et al., 2024] focuses on bi-temporal image analysis, supporting both change detection and change captioning, and enables interactive change interpretation based on user instructions.  [Zhu et al., 2024] proposes a LLM-guided multi-agent system for remote sensing image generation. RS-ChatGPT [Guo et al., 2024a] represents a more general-purpose attempt. By combining ChatGPT with a suite of pre-trained remote sensing image processing models, it can handle seven tasks such as object detection, scene classification, and image captioning in response to user queries. However, those agent models are limited by a relatively narrow toolkit, which restricts its ability to fully address complex, domain-specific requirements. Additionally, they often face challenges in flexible task planning and occasionally fail in reasoning or tool selection.
@@ 2.3 Retrieval-Augmented Generation
Large Language Models (LLM) excels in generating fluent text thanks to large-scale pre-training. However, their performance on domain-specific tasks—such as those in remote sensing—is limited by the lack of relevant or up-to-date training data. To mitigate this, Retrieval-Augmented Generation (RAG) has emerged as a promising approach, enabling language models to access external, domain-specific knowledge sources [Kenneweg et al., 2024].
RAG frameworks typically employ dense or sparse retrieval mechanisms to fetch relevant documents from a large corpus [Gao et al., 2023a, Gao et al., 2023b, Chan et al., 2024], which are then fused with the input prompt of LLM to guide the generation process. While effective for general use, they often fall short on complex queries requiring multi-hop reasoning or capturing comprehensive global information [Tang and Yang, 2024]. To improve this, knowledge graph-based RAG has been introduced [Edge et al., 2024, Guo et al., 2024b, Zhu et al., 2025], integrating structured entity-relation data into retrieval to support more accurate and semantically rich responses.
Among them, LightRAG [Guo et al., 2024b] proposed a lightweight framework that combines graph-based indexing with keyword-guided queries. It constructs a combined vector by concatenating generated keywords from query, capturing their semantic connections. While LightRAG introduces an efficient retrieval framework and leverages graph-based indexing, it suffers from two major drawbacks: (1) it treats all generated keywords equally, ignoring their varying importance in representing the query’s intent; and (2) it constructs a single concatenated query from all keywords, which may obscure the semantic distinctiveness of loosely related or heterogeneous concepts. This often results in reduced retrieval coverage and suboptimal relevance, particularly for complex, multi-aspect queries.
@@ 3.1 Overall Architecture
1:The user query $Q$ and the input remote sensing image $I$;
2:Answer $A$ generated by the RS-Agent; Processed remote sensing image $\hat{I}$ (if required by the task, i.e., object detection results, etc);
3:Central controller $M_{c}$ receives user query $Q$ and determines the task type $tp$;
4:Transmits $tp$ to solution searcher $M_{s}$;
5:$M_{s}$ searches through the Solution Database $D_{s}$, and retrieves the most relevant documents $d$ as Solution Guidance $g_{s}$;
@@
10:     Invoke $\hat{T}$ to process the image $I$, and get the corresponding outputs and $\hat{I}$;
@@
12:$M_{c}$ compiles the final answer $A$ from the output of $\hat{T}$;
13: return $A$ and $\hat{I}$;
@@ 3.2 Central Controller
The Central Controller $M_{c}$ serves as the brain of the RS-Agent to manage multi-step reasoning and tool execution through modular interfaces. We build $M_{c}$ using a large language model and leverage the LangChain framework [Chase, 2022] to support task decomposition, context management, and dynamic tool invocation. To improve task planning accuracy, the Central Controller $M_{c}$ adopts a structured, two-stage prompt-based reasoning process to guide precise task execution.
@@ 3.3 Toolkit
The Toolkit $T$ serves as the action engine of RS-Agent, comprising a collection of 18 tools tailored for specific remote sensing applications such as image caption [Wang et al., 2020a], scene classification [Xu et al., 2023], super resolution [Lin et al., 2020], etc. Details of the available tools are provided in Table 8 and illustrated in Figure 3. All tools are invoked through API-based interfaces, allowing for efficient and modular integration. This modular design ensures that the Toolkit can be continuously extended with additional tools, allowing RS-Agent to keep pace with emerging tasks, methods, and challenges in the remote sensing domain.
@@ 3.4 Solution Space
In the RS-Agent architecture, the Solution Space plays a critical role in supporting accurate task planning by compensating for the lack of domain-specific solution strategies in large language models (LLMs). While LLMs have shown impressive capabilities in natural language understanding and reasoning, they often struggle with remote sensing tasks that require specialized knowledge and precise procedural guidance. This limitation is especially evident when task instructions are ambiguous, underspecified, or implicitly formulated [Qin et al., 2023, Shapira et al., 2023]—scenarios common in real-world remote sensing applications.
@@ 3.4.2 Solution Searcher with Task-Aware Retrieval
Solution Retrieval: The inferred $tp$ is then passed to the $M_{s}$, which retrieves relevant strategies from $D_{s}$ to support tool selection and reasoning. To achieve the retrieval, each solution document in $D_{s}$ and the task type $tp$ are encoded into a dense vector representation with a sentence embedding model [Chen et al., 2024]. Then Solution Searcher $M_{s}$ employs the FAISS algorithm [Jégou et al., 2022], a high-performance vector similarity search algorithm $f_{s}$, to find the most relevant documents, and form the Solution Guidance $g_{s}=d$:
@@ 3.5.1 Knowledge Database
The Knowledge Database $D_{k}$ serves as a structured repository of domain knowledge that supports RS-Agent in tasks requiring factual accuracy, technical terminology, or background-specific reasoning. It provides a foundation for semantic retrieval and reasoning, allowing the agent to access reliable information when responding to complex or knowledge-intensive queries. In this paper, to theoretically validate RS-Agent’s capability in retrieving and generating domain-specific knowledge, we construct a small-scale knowledge database $D_{k}$ named RSaircraft for case-study. This curated dataset includes information on 65 types of aircraft, such as fighter jets, commercial airliners, drones, and helicopters. We adopt a knowledge graph-based storage structure [Guo et al., 2024b], in which each text chunk is processed by an LLM to extract entities and relational triples. These entities and relations are embedded into separate vector spaces, enabling structure-aware semantic retrieval. This representation supports multi-hop reasoning and domain-informed query expansion, facilitating more accurate and context-rich responses.
@@ 3.5.2 Knowledge Searcher with DualRAG
The retrieved documents form the Knowledge Guidance $g_{k}=d_{\text{global}}\cup d_{\text{ind}}$, which provides the LLM with domain-specific knowledge. By combining both retrieval paths, DualRAG achieves a better balance between semantic coherence and informational diversity, addressing the limitations of single-query strategies like that in LightRAG [Guo et al., 2024b].

### v3

@@ 1 Introduction
By aligning visual and textual information, Multi-modal Large Language Models (MLLMs) [41, 46] have made considerable progress across diverse remote sensing image interpretation tasks, including image captioning [49], visual question answering [2], and semantic scene understanding [22, 25], making remote sensing data more accessible to non-expert users. However, despite these advancements, MLLMs still face challenges in consistently handling multiple tasks across diverse remote sensing image modalities and resolutions. Firstly, remote sensing data encompass a wide range of modalities with varying resolutions. Consequently, general-purpose models lack the flexibility to quickly accommodate new modalities without computationally expensive retraining. In addition, most existing MLLMs are primarily optimized for instruction-following or descriptive tasks, such as image captioning and visual question answering. These generalist models typically do not provide expert-level responses comparable to those specialized models customized for specific tasks [47, 26, 58, 74]. These challenges emphasize the necessity for an expert-aware system capable of interacting with human users, and automatically utilizing specialized models to address the demands of real-world remote sensing applications.
It is believed that AI Agent [63] is a promising solution to these challenges. AI agents are autonomous systems capable of understanding user intent, planning multi-step tasks, invoking appropriate tools, and refining their decisions based on intermediate outcomes [42, 38, 57]. Recent advancements in agent frameworks—often powered by large language models—have demonstrated strong potential in coordinating multiple tools, decomposing complex tasks, and flexibly adapting to diverse user goals through modular design [43, 63]. These agents can bridge the gap between high-level semantic instructions and low-level tool execution operations. Despite growing interest, such agent-based approaches remain largely underexplored in the remote sensing domain. Although a few works in the remote sensing field use the term “agent” in their paper [12, 30, 73], they primarily focus on specific tasks and fail to propose a general-purpose agent. RS-ChatGPT [18] integrates ChatGPT [3] with a collection of pre-trained remote sensing networks to perform tasks such as object detection, scene classification, and image captioning. However, RS-ChatGPT encounters limitations when scaling to a broader range of tools or addressing tasks that demand specialized domain knowledge.
@@
The core functionality of RS-Agent lies in its ability to accurately interpret user instructions and convert them into precise and executable workflows. To this end, we design a carefully structured agent architecture that integrates modular system prompts with a long-term memory mechanism. This design empowers the Central Controller to accurately interpret user intent and dynamically plan workflows by leveraging current inputs, tool outputs, and contextual memory. Robustness and adaptability are critical for such systems, especially when handling diverse and ambiguous user instructions. Recent studies have emphasized the importance of stable task-level decision-making under varying task distributions and execution contexts [53, 35]. In the Solution Space, to empower RS-Agent with the ability to reason and operate like a professional remote sensing analyst, we build a Solution Database that stores professional knowledge on how to use various remote sensing tools effectively. We further propose a Task-Aware Retrieval method to retrieve and comprehend expert knowledge, enabling accurate tool selection and step-by-step task decomposition tailored to the user’s query. To address the domain knowledge limitations of general-purpose models [3, 50, 1], we develop a dedicated Knowledge Database, tailored to remote sensing targets. Building on this, we propose DualRAG, a novel retrieval-augmented generation method that enables the agent to retrieve more relevant and accurate domain-specific knowledge.
RS-Agent accurately interprets user intent and offers flexible integration of new tools to support emerging applications. We conduct extensive experiments across 9 datasets and 18 remote sensing tasks. Experiments reveal that the RS-Agent significantly enhances task planning accuracy—achieving over 95% accuracy, substantially outperforming the SOTA baseline RS-ChatGPT [18]. Furthermore, by coordinating multiple tools, RS-Agent consistently surpasses single-model performance across diverse applications, including scene classification, remote sensing visual question answering, and object counting. In addition, the proposed DualRAG method markedly improves the quality of domain-specific knowledge retrieval. It is also worth noting that this work does not aim to develop new LLM architectures. Instead, RS-Agent is designed to be compatible with both open-source and proprietary LLMs. We validate the generality and adaptability of our framework through experiments with several state-of-the-art models, including ChatGPT [3], LLaMA [50], Qwen [67], and DeepSeek [17], demonstrating the generality and adaptability of our framework.
@@ 2.1 Remote Sensing Multimodal Large Language Models
The integration of Multimodal Large Language Models (MLLMs) into remote sensing has opened new opportunities for human-AI interaction in earth observation tasks. MLLMs aim to bridge visual and semantic spaces, enabling the processing and interpretation of remote sensing imagery based on natural language inputs. Early applications of MLLMs in remote sensing primarily focused on a limited set of tasks, such as image captioning [45, 60] and visual question answering (VQA) [32]. As research progressed, MLLMs began to support a broader range of vision-language tasks [59], each addressing specific challenges in remote sensing interpretation. RSGPT [22] integrates both image captioning and VQA. GeoChat [25] extended this capability to visual geo-localization via interactive dialogue. LHRS-Bot [36] is a general-purpose MLLM built upon large-scale datasets and a novel multi-level vision-language alignment strategy, achieving state-of-the-art performance in classification, visual question answering, and visual grounding tasks. EarthVQANet [56] introduces a novel multi-task framework combining semantic segmentation and VQA to enhance spatial reasoning in remote sensing imagery. SkySenseGPT [33] leveraged knowledge graphs to enhance relational reasoning and complex target understanding. Beyond single-image interpretation tasks, a number of MLLM-based models have also been developed to address the interpretation of multi-scene remote sensing data. ChangeCLIP [10] filled a critical gap by tackling change captioning between multi-temporal images. SkyEyeGPT [70] introduced video captioning for remote sensing. UniRS [28] further advanced the field by integrating single-frame, bi-temporal, and video inputs into a unified model designed specifically for temporal remote sensing analysis.
@@ 2.2 AI Agent
AI agents are autonomous systems designed to perceive their environment, process information, make decisions, and execute actions to achieve specific objectives [62]. Equipped with the ability to learn from interactions, adapt to new data, and continuously improve performance, these agents are capable of functioning effectively across a wide range of application domains [38, 13].
While AI agents have been widely explored in general domains, their application in the field of remote sensing remains limited. Only a few agent-based systems have been proposed, typically addressing some specific defined tasks. TreeGPT [12], for instance, is designed for forestry applications, enabling individual tree segmentation and ecological parameter estimation from remote sensing imagery. Change-Agent [30] focuses on bi-temporal image analysis, supporting both change detection and change captioning, and enables interactive change interpretation based on user instructions. [73] proposes a LLM-guided multi-agent system for remote sensing image generation. RS-ChatGPT [18] represents a more general-purpose attempt. By combining ChatGPT with a suite of pre-trained remote sensing image processing models, it can handle seven tasks such as object detection, scene classification, and image captioning in response to user queries. However, those agent models are limited by a relatively narrow toolkit, which restricts its ability to fully address complex, domain-specific requirements. Additionally, they often face challenges in flexible task planning and occasionally fail in reasoning or tool selection.
@@ 2.3 Retrieval-Augmented Generation
Large Language Models (LLM) excels in generating fluent text thanks to large-scale pre-training. However, their performance on domain-specific tasks—such as those in remote sensing—is limited by the lack of relevant or up-to-date training data. To mitigate this, Retrieval-Augmented Generation (RAG) has emerged as a promising approach, enabling language models to access external, domain-specific knowledge sources [24].
RAG frameworks typically employ dense or sparse retrieval mechanisms to fetch relevant documents from a large corpus [15, 16, 4], which are then fused with the input prompt of LLM to guide the generation process. While effective for general use, they often fall short on complex queries requiring multi-hop reasoning or capturing comprehensive global information [48]. To improve this, knowledge graph-based RAG has been introduced [14, 19, 75], integrating structured entity-relation data into retrieval to support more accurate and semantically rich responses.
Among them, LightRAG [19] proposed a lightweight framework that combines graph-based indexing with keyword-guided queries. It constructs a combined vector by concatenating generated keywords from query, capturing their semantic connections. While LightRAG introduces an efficient retrieval framework and leverages graph-based indexing, it suffers from two major drawbacks: (1) it treats all generated keywords equally, ignoring their varying importance in representing the query’s intent; and (2) it constructs a single concatenated query from all keywords, which may obscure the semantic distinctiveness of loosely related or heterogeneous concepts. This often results in reduced retrieval coverage and suboptimal relevance, particularly for complex, multi-aspect queries.
@@ 3.1 Overall Architecture
1:The user query $Q$ and the input remote sensing image II;
2:Answer AA generated by the RS-Agent; Processed remote sensing image $\hat{I}$ (if required by the task, i.e., object detection results, etc);
3:Central controller $M_{c}$ receives user query $Q$ and determines the task type tptp;
4:Transmits tptp to solution searcher $M_{s}$;
[…]
@@
10:  Invoke $\hat{T}$ to process the image II, and get the corresponding outputs and $\hat{I}$;
@@
[…]
13: return AA and $\hat{I}$;
@@ 3.2 Central Controller
The Central Controller $M_{c}$ serves as the brain of the RS-Agent to manage multi-step reasoning and tool execution through modular interfaces. We build $M_{c}$ using a large language model and leverage the LangChain framework [5] to support task decomposition, context management, and dynamic tool invocation. To improve task planning accuracy, the Central Controller $M_{c}$ adopts a structured, two-stage prompt-based reasoning process to guide precise task execution.
@@ 3.3 Toolkit
The Toolkit $T$ serves as the action engine of RS-Agent, comprising a collection of 18 tools tailored for specific remote sensing applications such as image caption [55], scene classification [66], super resolution [29], etc. Details of the available tools are provided in Table 8 and illustrated in Figure 3. All tools are invoked through API-based interfaces, allowing for efficient and modular integration. This modular design ensures that the Toolkit can be continuously extended with additional tools, allowing RS-Agent to keep pace with emerging tasks, methods, and challenges in the remote sensing domain.
@@ 3.4 Solution Space
In the RS-Agent architecture, the Solution Space plays a critical role in supporting accurate task planning by compensating for the lack of domain-specific solution strategies in large language models (LLMs). While LLMs have shown impressive capabilities in natural language understanding and reasoning, they often struggle with remote sensing tasks that require specialized knowledge and precise procedural guidance. This limitation is especially evident when task instructions are ambiguous, underspecified, or implicitly formulated [40, 44]—scenarios common in real-world remote sensing applications.
@@ 3.4.2 Solution Searcher with Task-Aware Retrieval
Solution Retrieval: The inferred $tp$ is then passed to the $M_{s}$, which retrieves relevant strategies from $D_{s}$ to support tool selection and reasoning. To achieve the retrieval, each solution document in $D_{s}$ and the task type $tp$ are encoded into a dense vector representation with a sentence embedding model [6]. Then Solution Searcher $M_{s}$ employs the FAISS algorithm [23], a high-performance vector similarity search algorithm $f_{s}$, to find the most relevant documents, and form the Solution Guidance $g_{s}=d$:
@@ 3.5.1 Knowledge Database
The Knowledge Database $D_{k}$ serves as a structured repository of domain knowledge that supports RS-Agent in tasks requiring factual accuracy, technical terminology, or background-specific reasoning. It provides a foundation for semantic retrieval and reasoning, allowing the agent to access reliable information when responding to complex or knowledge-intensive queries. In this paper, to theoretically validate RS-Agent’s capability in retrieving and generating domain-specific knowledge, we construct a small-scale knowledge database $D_{k}$ named RSaircraft for case-study. This curated dataset includes information on 65 types of aircraft, such as fighter jets, commercial airliners, drones, and helicopters. We adopt a knowledge graph-based storage structure [19], in which each text chunk is processed by an LLM to extract entities and relational triples. These entities and relations are embedded into separate vector spaces, enabling structure-aware semantic retrieval. This representation supports multi-hop reasoning and domain-informed query expansion, facilitating more accurate and context-rich responses.
@@ 3.5.2 Knowledge Searcher with DualRAG
The retrieved documents form the Knowledge Guidance $g_{k}=d_{\text{global}}\cup d_{\text{ind}}$, which provides the LLM with domain-specific knowledge. By combining both retrieval paths, DualRAG achieves a better balance between semantic coherence and informational diversity, addressing the limitations of single-query strategies like that in LightRAG [19].